OpenStudy (anonymous):

4x - 5y = 7 -4x + 5y = 7 How do I solve this equation using the addition method? I'm stuck as 4x and -4x cancel out, so do -5y and 5y which leaves me with 0 = 14...

Directrix (directrix):

If this is a system of equations for which you are looking for a common solution: 4x - 5y = 7 -4x + 5y = 7 ------------ 0 = 14 indicates that the two lines do not intersect.

Directrix (directrix):

Good site for review: http://www.themathpage.com/alg/simultaneous-equations.htm